# A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status

### v3

@@ 5.2. Limitations
Our principal aim has been to illustrate how prior-infection protection, as well as comorbidity status, can be integrated into benefit-risk modeling of a two-dose vaccination series. In applying our model to extend and improve the FDA’s mRNA-1273 assessment [11] relative to its hospitalizations endpoint, we have assumed that young males are hospitalized for COVID at a lower rate than young females, as did the FDA in [10] and [11]. Data from [10] and [54] suggests this rate difference is substantial; however, the rates of [10] and [54] include incidental hospitalizations and incidental rates for males may differ from those for females. We have assumed, as did the FDA, that vaccination and prior infection offer essentially equivalent protection for males and females. Also, we have based our assessment of the difference in hospitalization rates between young males and females on data in Table 2 of [10] representing mostly pre-Delta cases. Relative to Omicron, sex-based differences in hospitalization rates might have changed. However, the rate differences provided by Table 1 of [54], based on two years of data including that of the initial Omicron wave, are consistent with those of Table 2 of [10].
@@
COVID-hospitalizations data from Ontario [Supplement 2] and from Connecticut [Supplement 1, Appendix 1] suggests that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25 year-olds (general population mix of males and females). Thus, even if we assume males and females in this age range are hospitalized for COVID at roughly equivalent rates, our IHR value of 0.14% for 18–25-year-old males appears to be a good estimate.
@@
In Section 2 of Supplement 1, based on data available before 1/22/22, we obtain an Omicron IHR estimate for 18–25 year-olds (general-population mix of males and females) of 0.22%, not 0.14%. One possible explanation is that we may have underestimated the incidental COVID-hospitalization rate for 18–25 year-olds. Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental Omicron-hospitalization rate of 40%, a rate suggested by [28] and [30] for hospitalizations with COVID among patients of all ages. However, as noted in Section 2.2 (and discussed in detail in Supplement 1), there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly. In particular, COVID-hospitalization data discussed in Appendix 1 of Supplement 1 suggests an incidental rate of over 70% among 15–24 year-olds in Connecticut during calendar year 2020. Assuming an incidental rate of about 62.2% (instead of 40%) transforms our IHR estimate of 0.22% into the estimate 0.14%.
@@
Our risk assessment, like the FDA’s, ignores known, rare vaccine-related adverse events other than myo/pericarditis that may require hospitalization. Finally, our risk assessment, like the FDA’s, ignores the possibility that COVID-19 vaccination may have long-term health effects; for instance, the FDA’s package-insert information for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals” [78, Section 13].

### v4

@@ 4.3. Limitations
Our principal aim has been to illustrate how prior-infection protection, as well as comorbidity status, can be integrated into a deterministic benefit–risk model that extends, in a natural way, the one the FDA employed in its assessment of mRNA-1273 vaccination. We have developed and presented our model in accordance with BRIVAC (Benefit–RIsk models applied to VACcines) guidelines, in particular, its item 17 “Sensitivity/scenario analyses” [80]. We acknowledge that it is desirable to employ more sophisticated approaches to benefit–risk modeling and to accounting for uncertainty in such modeling; for instance, Monte Carlo simulation modeling with Beta/Poisson distributions for model inputs, guiding input choices for simulations.
@@
In applying our model to extend and improve the FDA’s mRNA-1273 assessment [13] relative to its hospitalizations endpoint, we have assumed that young males are hospitalized for COVID at a lower rate than young females, as did the FDA in [12, 13]. Data from [12, 57] suggests this rate difference is substantial; however, the rates of [12, 57] include incidental hospitalizations, and incidental rates for males may differ from those for females. We have assumed, as did the FDA, that vaccination and prior infection offer essentially equivalent protection for males and females. Also, we have based our assessment of the difference in hospitalization rates between young males and females on data in Table 2 of [12] representing mostly pre-Delta cases. Relative to Omicron, sex-based differences in hospitalization rates might have changed. However, the rate differences provided by Table 1 of [57], based on two years of data including that of the initial Omicron wave, are consistent with those of Table 2 of [12].
@@
COVID-hospitalization data from Ontario [Supplement S2] and from Connecticut [Supplement S1, Appendix S1] suggests that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25-year-olds (general population mix of males and females). Thus, even if we assume males and females in this age range are hospitalized for COVID at roughly equivalent rates, our IHR value of 0.14% for 18–25-year-old males appears to be a good estimate.
@@
In Section S2 of Supplement S1, based on data available before 22 January 2022, we obtain an Omicron IHR estimate for 18–25 year-olds (general-population mix of males and females) of 0.22%, not 0.14%. One possible explanation is that we may have underestimated the incidental COVID-hospitalization rate for 18–25 year-olds. Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental Omicron-hospitalization rate of 40%, a rate suggested by [34, 36] for hospitalizations with COVID among patients of all ages. However, as noted in Section 1.3 (and discussed in detail in Supplement S1), there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly. In particular, COVID-hospitalization data discussed in Appendix S1 of Supplement S1 suggests an incidental rate of over 70% among 15–24-year-olds in Connecticut during calendar year 2020. Assuming an incidental rate of about 62.2% (instead of 40%) transforms our IHR estimate of 0.22% into the estimate 0.14%.
@@
Our risk assessment, like the FDA’s, ignores known, rare vaccine-related adverse events other than myo/pericarditis that may require hospitalization. Finally, our risk assessment, like the FDA’s, ignores the possibility that COVID-19 vaccination may have long-term health effects; for instance, the FDA’s package-insert information for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals” [81, Section 13].
5. Conclusions
Our model illustrates the importance of nuanced vaccine recommendations based on careful benefit–risk modeling that takes into account health status and prior infection(s), two covariates that were not included in the FDA’s models [12, 13]. For healthy young males with prior infection, our model suggests that risks outweighed benefits (relative to hospitalizations) based on data available prior to licensure of the mRNA-1273 product. However, for higher-risk young males with obesity, and especially those with no prior infection, vaccination was found to confer benefit.
Risks of infection and vaccination, such as long COVID and VAM/P, must be carefully weighed to offer informed consent. Current long-COVID prevalence estimates among adults with confirmed SARS-CoV-2 infection are most reliably estimated in the range of 7–15% when using standardized definitions and representative sampling [82, 83], and when using a conservative definition consistent with other post-viral syndromes, the WHO finds 2.8% among those younger than 20 [84].
With respect to VAM/P, more recent formulations of Moderna’s COVID vaccine have a reduced mRNA dose and VAM/P-incidence rates have declined [85]. Long-term VAM/P prognosis remains somewhat unclear. A CDC follow-up at 90 days found that although the majority (81%) of VAM/P cases recovered, 54% had ongoing abnormal cardiac findings and 26% were on daily cardiac medication; only 68% were cleared for all physical activity [86]. Clinical guidelines for return-to-sport following myocarditis call for 3–6 months’ activity restriction [87, 88]. In addition, there is the unknown prognosis for complications like late gadolinium enhancement (LGE) [89].
The question of boosting the general population, including healthy young males, is beyond the scope of this paper and requires tremendous nuance, a fact widely recognized by fall 2021 [90]. The FDA indicated in May 2025 [91] that it anticipates needing data from RCTs to evaluate clinical outcomes prior to authorizing COVID-19 vaccines for healthy individuals ages 6 months to 64 years, while immunogenicity standards will likely suffice for those with at least one chronic condition and those over age 65. The agency’s move thus operationalizes the risk-stratified approach supported by our deterministic modeling and will generate data that can be used in more sophisticated modeling.